First LBC tech Night. LBC team present.
Nautical Twilights: 01:42 UT 12:28 UT

01:42 Closed due to strong winds and high RH

04:05 Weather still bad with RH~88% and Wind up to 26m/s. LBC Team decides to dismount LBCB to inspect one of the two motors of the shutter curtains that does not respond

05:00 Still closed due to bad weather; ONE of the motors has been found to have been damaged, possibly during the LBCB re-mounting.

10:00 Weather still bad. Call it a night.

-- MarcoPedani - 13 Oct 2009
I Attachment Action Size Date Who Comment
daq.20091013.gifgif daq.20091013.gif manage 29 K 14 Oct 2009 - 01:23 JohnHill Plot of Telescope and Mirror Temperatures
Topic revision: r3 - 14 Oct 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback